mikewwe13:

The total amount of money in a savings account after t years is given by the function A=1000(1.023)t . How could this function be rewritten to identify the monthly interest rate? What is the approximate monthly interest rate? Drag and drop the choices into the boxes to correctly complete the table. If a value does not match, do not drag it to the table. Function Monthly interest rate A=1000(1+0.023) ^12t A=1000 (1.0231^12 )^12t A=1000 (1.023^12)^t12 0.19% 0.23% 0.31%

mikewwe13:

@Vocaloid

Vocaloid:

To find the monthly interest rate divide the yearly interest rate by 12

Vocaloid:

Since the yearly interest rate is 2.3% it's just 2.3%/12

Vocaloid:

However I am not sure about which function you are supposed to pick b/c none of those look quite right to me

mikewwe13:

let me fix it for you

mikewwe13:

The total amount of money in a savings account after t years is given by the function A=1000(1.023)t . How could this function be rewritten to identify the monthly interest rate? What is the approximate monthly interest rate? Drag and drop the choices into the boxes to correctly complete the table. If a value does not match, do not drag it to the table. Function Monthly interest rate A=1000(1+0.023) ^12t A=1000 (1.0231^12 )^12t A=1000 (1.023^12)^t12 0.19% 0.23% 0.31%

mikewwe13:

does this help ?

Vocaloid:

I got nothing @sillybilly123 halp me

mikewwe13:

The initial number of views for a blog was 20. The number of views is growing exponentially at a rate of 20% per week. What is the number of views expected to be four weeks from now? Round to the nearest whole number. Enter your answer in the box. __________

sillybilly123:

dividing the annual rate by 12 is a great approx and nothing wrong. but **if** they're **compounding on a monthly basis**, the norm, then you can go further into the detail If compounding interest on monthly basis at monthly rate \(i_m\), we have: \((1 + 0.023) = (1+i_m)^{12} \qquad \star\) \(\implies (1 + 0.023)^t = ((1+i_m)^{12})^ t\), where t connotes years \(= (1+i_m)^ {12 t}\) **but** go back to \(\star\) to work out the monthly compounded rate, \(i_m\), or just play around with the series.